# Decoupling and Damping: Structurally-Regularized Gradient Matching for Multimodal Graph Condensation

### v1

@@ Abstract.
In critical web applications such as e-commerce and recommendation systems, multimodal graphs integrating rich visual and textual attributes are increasingly central, yet their large scale introduces substantial computational burdens for training Graph Neural Networks (GNNs). While Graph Condensation (GC) offers a promising solution by synthesizing smaller datasets, existing methods falter in the multimodal setting. We identify a dual challenge causing this failure: (1) conflicting gradients arising from semantic misalignments between modalities, and (2) the GNN’s message-passing architecture pathologically amplifying this gradient noise across the graph structure. To address this, we propose Structurally-Regularized Gradient Matching (SR-GM), a novel condensation framework tailored for multimodal graphs. SR-GM introduces two synergistic components: first, a gradient decoupling mechanism that resolves inter-modality conflicts at their source via orthogonal projection; and second, a structural damping regularizer that acts directly on the gradient field. By leveraging the graph’s Dirichlet energy, this regularizer transforms the topology from a noise amplifier into a stabilizing force during optimization. Extensive experiments demonstrate that SR-GM significantly improves accuracy and accelerates convergence compared to baseline methods. Ablation studies confirm that addressing both gradient conflict and structural amplification in tandem is essential for achieving superior performance. Moreover, the condensed multimodal graphs exhibit strong cross-architecture generalization and promise to accelerate applications like Neural Architecture Search. This research provides a scalable methodology for multimodal graph-based learning in resource-constrained environments.
@@ 1. Introduction
The deep integration of the Internet and multimedia technologies has spawned a massive amount of graph-structured data containing rich heterogeneous information. From users in social networks and the multimedia content they publish (khattar2019mvae; wang2012multimodal; su2021detecting; tao2020mgat; safavi2020codex) to product graphs associated with text descriptions and product images on e-commerce platforms (xu2020product; deng2023construction), these multimodal graphs have become the core of modern web applications. Graph neural networks (GNNs) (kipf2016semi; rong2019dropedge; zhang2018link), as powerful tools for processing such data, have achieved remarkable results in tasks such as node classification and link prediction. However, their success relies heavily on large-scale datasets (safavi2020codex; wan2023scalable), which leads to high computational costs, including lengthy training time, huge memory consumption, and complex hyperparameter tuning (zhang2022efficient; cai2024multimodal).
@@
To address this challenge, a data-centric approach, dataset condensation (wang2018dataset; zhao2020dataset), has emerged. Its goal is to refine the original large-scale dataset into a small synthetic dataset with highly concentrated information, so that the model trained on this synthetic dataset can achieve performance comparable to that trained on the full dataset. In the field of computer vision, gradient matching (zhao2020dataset) has become a benchmark method for achieving this goal. It efficiently learns highly representative synthetic samples by matching the model’s training gradient trajectories on real data and synthetic data. Recently, researchers have begun to extend this paradigm to graph-structured data and proposed a variety of graph condensation methods (jin2021graph; gao2023multiple; liu2022graph). These methods are usually formulated as a complex bi-level optimization problem, aiming to learn a miniature synthetic graph structure and its node features through objectives such as gradient matching (jin2021graph) or distribution matching (liu2022graph).
@@
Although these methods show potential on unimodal graphs (with node features from a unimodal source), their performance degrades significantly when directly applied to more general multimodal graphs (where nodes have multiple modal features such as text and images). Through in-depth analysis, we find that the root cause lies in the conflict between multimodal gradient signals and the pathological amplification effect of the graph structure on this conflict. Specifically, there may be a semantic gap between different modalities (such as text and image) (liang2022mind; zhang2025beyond), which leads to the generation of gradients in opposite directions during the optimization process. More importantly, the inherent message passing mechanism of GNN will quickly propagate and amplify the gradient conflicts of these local nodes to the entire graph, thus making the optimization process of the condensation algorithm based on gradient matching unstable and ineffective.
@@
To overcome this problem, this paper proposes a novel structurally regularized gradient matching (SR-GM) framework designed for multimodal graph condensation. The core ideas of SR-GM are twofold: (1) Gradient decoupling: At its core, we adopt the idea of conflict disentanglement from multi-task learning (yu2020gradient; chai2024towards). Specifically, for the conflicting gradients arising from different modalities on a synthetic node, we apply orthogonal projection to eliminate their inherent contradictions, thereby achieving synergistic inter-modal gradient collaboration; (2) Structural damping: On the propagation path, we innovatively introduce a graph Laplacian regularization term that acts on the gradient field rather than node features. This regularization term forces the optimization directions of adjacent nodes to be consistent during the condensation process, thereby transforming the graph structure from a ”conflict amplifier” to an ”optimization damper”, effectively suppressing the spread of gradient noise.
@@
We conducted extensive experiments on multiple multimodal graph benchmark datasets. Our results demonstrate that SR-GM significantly outperforms existing graph condensation baselines. Detailed ablation studies confirm the necessity and effectiveness of the two components, gradient decoupling and structural damping. Furthermore, the condensed graphs generated by SR-GM exhibit excellent generalization across diverse GNN architectures, highlighting its robustness and practicality.
@@
2. Related Work
2.1. Graph Dataset Condensation
Dataset Condensation (DC), also known as Dataset Distillation, aims to synthesize a small, informative dataset from a large original one such that models trained on the synthetic set achieve performance comparable to those trained on the original data. This paradigm was first pioneered in the image domain, where methods were developed to optimize synthetic pixels directly through a bi-level optimization framework (wang2018dataset). The core idea is to match certain properties between the real and synthetic data. Various matching criteria have been proposed, including matching model performance (zhao2020dataset) and feature distributions (zhao2023dataset; wang2022cafe). Among these, matching the gradients of model parameters has been proven to be a particularly effective strategy (zhao2020dataset). The extension of these techniques to the graph domain, termed Graph Condensation (GC), is a nascent and challenging endeavor. Jin et al. (jin2021graph) first extended DC to graphs, introducing a node-level condensation method. Unlike image data, graph data involves an intricate coupling between node attributes and topological structure, making the joint synthesis of features and edges a complex optimization task. Current GC methods primarily focus on preserving information in single-modal graphs, exploring strategies such as gradient matching (jin2021graph), training trajectory matching (zheng2023structure), and distribution matching (liu2022graph). Recent research has evolved from developing efficient and interpretable graph condensation methods (fang2024exgc; wang2024fast) to creating specialized techniques tailored for specific graph types, such as heterogeneous graphs and signed graphs (gao2024heterogeneous; li2025structure). However, these methods largely overlook the unique challenges posed by multimodal graphs, where node features themselves are composed of heterogeneous information from sources like text and images. The problem of inter-modal signal conflict within the GC framework remains a critical and underexplored gap. Our work aims to bridge this gap.
2.2. Multimodal Graph Learning
Multimodal Graph Learning (MGL) has emerged as a vital field for analyzing complex real-world systems where entities are described by multiple data types (e.g., text, images, audio). MGL aims to leverage the relational structure of graphs to fully explore associations both within and between different modalities. This is particularly challenging because data fusion must be guided by the graph topology, which itself may vary across modalities. A key challenge in MGL is bridging the ”semantic gap”, where different modalities may provide conflicting or inconsistent representations of the same entity (zhu2025mosaic; liang2022mind). For instance, two products might be visually similar but have different textual descriptions, creating discrepancies that unimodal models cannot resolve. To address this, significant research has focused on feature alignment, which aims to map representations from different modalities into a unified embedding space. Contrastive learning frameworks, such as CLIP (radford2021learning), have been notably successful in learning aligned vision-language representations. From an architectural perspective, methods like Multimodal Graph Convolutional Networks (MGCN) (wei2019mmgcn) and Multimodal Graph Attention Networks (MGAT) (tao2020mgat) have been proposed to explicitly model and fuse information from different modalities. While these MGL methods focus on designing more sophisticated Graph Neural Network (GNN) architectures to better handle existing multimodal data, they do not address the data-centric challenge of scalability (ninggraph4mm; cai2024multimodal). The question of how to effectively compress large-scale multimodal graphs, particularly in a way that handles the feature conflicts inherent in the condensation process, remains largely unaddressed. Our work bridges this gap by introducing a condensation method specifically designed with an optimization scheme attuned to the uniqueness of multimodal graphs.
[…]
@@ 3.2. Gradient Matching for Multimodal Graph
This goal is formally captured as a bi-level optimization problem (jin2022condensing). The outer loop seeks to find the optimal synthetic graph $\mathcal{S}^{*}$ by minimizing a loss $\mathcal{L}_{\text{outer}}$ that evaluates a model trained on $\mathcal{S}$. The inner loop trains this model on $\mathcal{S}$ to convergence to obtain the optimal parameters $\theta_{\mathcal{S}}^{*}$. This can be written as:
@@
Solving this formulation directly is computationally prohibitive, as it requires fully training a GNN in the inner loop for every single update to the synthetic graph $\mathcal{S}$ in the outer loop.
The Gradient Matching (GM) paradigm offers an efficient and effective approximation to this problem (jin2021graph). Instead of matching the performance of fully converged models, GM aims to align the learning dynamics by matching the parameter gradients at each step of a single training trajectory. The process involves two interleaved optimizations: an outer-loop optimization for the synthetic data $\mathcal{S}$ and an inner-loop update for the model parameters $\theta$. The objective is to minimize the cumulative distance between the gradients over the entire trajectory:
@@
where the model parameters are updated in an inner loop according to the synthetic data: $\theta_{t+1}=\theta_{t}-\eta_{\theta}\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t})$. Here, $\mathcal{L}_{\mathcal{S}}$ and $\mathcal{L}_{\mathcal{T}}$ are the task losses, and $D(\cdot,\cdot)$ is a distance metric between the gradient vectors.
Traditional gradient matching methods, while successful in unimodal domains, face significant challenges when applied to multimodal graphs due to two fundamental issues.
@@ 3.3. Inter-Modality Gradient Conflict
The outer loop of the GM algorithm is tasked with updating the synthetic graph $\mathcal{S}$ to minimize the distance to a target gradient, $\mathbf{g}_{\mathcal{T}}=\nabla_{\theta}\mathcal{L}_{\mathcal{T}}$, sampled from the real data. In the multimodal setting, this target gradient becomes inherently unreliable.
To formalize this, we frame the GNN’s learning process as a Multi-Task Learning (MTL) problem (liu2021conflict; wang2020gradient), where each modality constitutes a distinct task being learned by the shared backbone, parameterized by $\theta_{\text{sh}}$. The total gradient $\mathbf{g}_{\mathcal{T}}$ is a linear combination of the gradients from each modality:
(3) $\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{total}}=\sum_{k\in\{\text{text, image}\}}w_{k}\nabla_{\theta_{\text{sh}}}\mathcal{L}_{k}$
This simple summation becomes problematic due to the feature-level ”semantic gap” inherent in multimodal data (feng2023mkvse; liu2024rng). This misalignment generates opposing update signals, leading to a gradient conflict(mahapatra2020multi; ma2020efficient), formally defined as the condition where the gradients of two tasks have a negative cosine similarity:
(4) $\langle\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{text}},\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{image}}\rangle<0$
When this occurs, the target gradient $\mathbf{g}_{\mathcal{T}}$ becomes a high-variance and suboptimal signal. It represents a blind compromise between conflicting objectives rather than a stable, Pareto-optimal update direction (mahapatra2020multi). Consequently, the outer loop of GM is tasked with matching a noisy, erratically moving target, which fundamentally destabilizes the meta-optimization of $\mathcal{S}$.
@@ 3.4. Structural Amplification of Gradient Noise
The instability is further compounded by the GNN’s message-passing mechanism (feng2022powerful), which corrupts the update step of the outer loop itself. The outer loop updates the synthetic data $\mathcal{S}$ using a meta-gradient, computed by differentiating the matching loss with respect to $\mathcal{S}$. This requires backpropagating through the computation of the synthetic gradient $\mathbf{g}_{\mathcal{S}}$, and therefore, through the GNN architecture.
@@
During this meta-gradient computation, the gradient of the loss with respect to a node’s features is influenced by the gradients of its neighbors, causing the gradient to flow backward along the graph edges:
(5) $\nabla_{h_{i}^{(k-1)}}\mathcal{L}=\sum_{j\mid i\in\mathcal{N}(j)}\frac{\partial\mathcal{L}}{\partial h_{j}^{(k)}}\frac{\partial h_{j}^{(k)}}{\partial h_{i}^{(k-1)}}$
@@
This dynamic transforms the GNN’s computational graph into a medium for noise propagation. The localized gradient conflicts at individual nodes are not contained; instead, they are amplified and propagated throughout the graph structure. This creates a highly non-smooth gradient field over the synthetic graph, where the optimization directions for adjacent nodes can be wildly different.
This structural amplification can be formally quantified using the Dirichlet Energy (zhou2021dirichlet) of the feature gradient field. Let $\mathbf{G}_{\mathcal{S}}$ be the $N^{\prime}\times d$ matrix where each row is the feature gradient $\nabla_{\mathbf{X}_{\mathcal{S},i}}\mathcal{L}$ for a synthetic node $i$, and let $\mathbf{L}_{\mathcal{S}}$ be the graph Laplacian. The Dirichlet energy of the gradient field is:
(6) $E(\mathbf{G}_{\mathcal{S}})$ $=\mathrm{tr}(\mathbf{G}_{\mathcal{S}}^{T}\mathbf{L}_{\mathcal{S}}\mathbf{G}_{\mathcal{S}})$
(7) $=\frac{1}{2}\sum_{(i,j)\in E_{\mathcal{S}}}w_{ij}\|\nabla_{\mathbf{X}_{\mathcal{S},i}}\mathcal{L}-\nabla_{\mathbf{X}_{\mathcal{S},j}}\mathcal{L}\|_{2}^{2}$
A high Dirichlet energy signifies a chaotic gradient field. The structural amplification mechanism ensures this energy remains high, creating a rugged optimization landscape for $\mathcal{S}$ and causing the meta-gradient to be unstable. This persistent instability in the gradient field directly undermines the condensation objective, making it difficult to learn a small yet representative synthetic graph.

### v2

@@ Abstract.
In multimodal graph learning, graph structures that integrate information from multiple sources, such as vision and text, can more comprehensively model complex entity relationships. However, the continuous growth of their data scale poses a significant computational bottleneck for training. Graph condensation methods provide a feasible path forward by synthesizing compact and representative datasets. Nevertheless, existing condensation approaches generally suffer from performance limitations in multimodal scenarios, mainly due to two reasons: (1) semantic misalignment between different modalities leads to gradient conflicts; (2) the message-passing mechanism of graph neural networks further structurally amplifies such gradient noise. Based on this, we propose Structural Regularized Gradient Matching (SR-GM), a condensation framework for multimodal graphs. This method alleviates gradient conflicts between modalities through a gradient decoupling mechanism and introduces a structural damping regularizer to suppress the propagation of gradient noise in the topology, thereby transforming the graph structure from a noise amplifier into a training stabilizer. Extensive experiments on four multimodal graph datasets demonstrate the effectiveness of SR-GM, highlighting its state-of-the-art performance and cross-architecture generalization capabilities in multimodal graph dataset condensation.
@@ 1. Introduction
The deep integration of the Internet and multimedia technologies has spawned a massive amount of graph-structured data containing rich heterogeneous information. From users in social networks and the multimedia content they publish (Khattar et al., 2019; Wang et al., 2012; Su et al., 2021; Tao et al., 2020; Safavi and Koutra, 2020) to product graphs associated with text descriptions and product images on e-commerce platforms (Xu et al., 2020; Deng et al., 2023), these multimodal graphs have become the core of modern web applications. Graph neural networks (GNNs) (Kipf, 2016; Rong et al., 2019; Zhang and Chen, 2018), as powerful tools for processing such data, have achieved remarkable results in tasks such as node classification and link prediction. However, their success relies heavily on large-scale datasets (Safavi and Koutra, 2020; Wan et al., 2023), which leads to high computational costs, including lengthy training time, huge memory consumption, and complex hyperparameter tuning (Zhang et al., 2022; Cai et al., 2024).
@@
To address this challenge, a data-centric approach, dataset condensation (Wang et al., 2018; Zhao et al., 2020), has emerged. Its goal is to refine the original large-scale dataset into a small synthetic dataset with highly concentrated information, so that the model trained on this synthetic dataset can achieve performance comparable to that trained on the full dataset. In the field of computer vision, gradient matching (Zhao et al., 2020) has become a benchmark method for achieving this goal. It efficiently learns highly representative synthetic samples by matching the model’s training gradient trajectories on real data and synthetic data. Recently, researchers have begun to extend this paradigm to graph-structured data and proposed a variety of graph condensation methods (Jin et al., 2021; Gao and Wu, 2023; Liu et al., 2022). These methods are usually formulated as a complex bi-level optimization problem, aiming to learn a miniature synthetic graph structure and its node features through objectives such as gradient matching (Jin et al., 2021) or distribution matching (Liu et al., 2022).
@@
Although these methods show potential on unimodal graphs where node features come from a unimodal source, their performance degrades significantly when directly applied to more general multimodal graphs where nodes have multiple modal features, such as text and images. Through in-depth analysis, we find that the root cause lies in the conflict between multimodal gradient signals and the pathological amplification effect of the graph structure on this conflict. Specifically, there may be a semantic gap between different modalities (Liang et al., 2022; Zhang et al., 2025), which leads to the generation of gradients in opposite directions during the optimization process. More importantly, the inherent message passing mechanism of GNN will quickly propagate and amplify the gradient conflicts of these local nodes to the entire graph, thus making the optimization process of the condensation algorithm based on gradient matching unstable and ineffective.
@@
To overcome this problem, we propose a novel structurally regularized gradient matching (SR-GM) framework designed for multimodal graph condensation. The core ideas of SR-GM are twofold: (1) Gradient decoupling: We adopt the idea of conflict disentanglement from multi-task learning (Yu et al., 2020; Chai et al., 2024). Specifically, for the conflicting gradients arising from different modalities on a synthetic node, we apply orthogonal projection to eliminate their inherent contradictions, thereby achieving synergistic inter-modal gradient collaboration; (2) Structural damping: On the propagation path, we innovatively introduce a graph Laplacian regularization term that acts on the gradient field rather than node features. This regularization term forces the optimization directions of adjacent nodes to be consistent during the condensation process, thereby transforming the graph structure from a ”conflict amplifier” to an ”optimization damper”, effectively suppressing the spread of gradient noise.
@@
We conducted extensive experiments on multiple multimodal graph benchmark datasets. Our results demonstrate that SR-GM significantly outperforms existing graph condensation baselines. Detailed ablation studies confirm the necessity and effectiveness of the two components, gradient decoupling and structural damping. Furthermore, the condensed graphs generated by SR-GM exhibit excellent generalization across diverse GNN architectures, highlighting their robustness and practicality.
@@
Figure 1. (a): Gradient matching for multimodal Graph. (b): Intra-modal similarity trends during condensation. Gradient conflict and noise amplification cause modality collapse. In this process, text features move closer to visual features, which raises the intra-modal cosine similarity (blue curve). Features then cluster more tightly, driving the concentration ratio (CR) [^1] A higher CR (Zhang et al., 2025) indicates that the given similarity corresponds to a narrower directional region on the hypersphere, implying stronger feature concentration in the high-dimensional embedding space.higher.
Figure 2. This is the overall framework of SR-GM. $\mathcal{T}$ and $\mathcal{S}$ denote the target and condensed multimodal graph datasets, respectively. $\theta_{t}(\mathcal{S})$ denotes the parameters of the GNN trained on the synthetic graph $\mathcal{S}$, while $\nabla_{\mathbf{X}_{\mathcal{S}}}\ell_{\mathcal{S}}(\theta_{t}(\mathcal{S}))$ denotes the gradient with respect to the synthetic node features $\mathbf{X}_{\mathcal{S}}$.
@@ 2.2. Gradient Matching for Multimodal Graph
To address the problem of multimodal graph dataset condensation, we adopt the graph dataset condensation paradigm. Specifically, the objective is formalized as a bi-level optimization problem (Jin et al., 2022). The outer loop aims to find the optimal synthetic graph $\mathcal{S}^{*}$ by minimizing a validation loss $\mathcal{L}_{\text{outer}}$ that evaluates a model trained on $\mathcal{S}$, while the inner loop trains the model on $\mathcal{S}$ to convergence to obtain the optimal parameters $\theta_{\mathcal{S}}^{*}$:
@@
Directly solving this bi-level optimization is computationally prohibitive, as each update to the synthetic graph $\mathcal{S}$ in the outer loop requires a complete inner-loop training of a GNN.
The Gradient Matching (GM) paradigm provides an effective approximation to this problem (Jin et al., 2021). Rather than matching the performance of fully converged models, GM aligns the learning dynamics by matching the parameter gradients at each step along a single training trajectory. The process consists of two interleaved optimizations: an outer-loop update for the synthetic data $\mathcal{S}$ and an inner-loop update for the model parameters $\theta$. The objective is to minimize the cumulative distance between the gradients over the trajectory:
@@
Here, model parameters are updated in the inner loop using the synthetic data: $\theta_{t+1}=\theta_{t}-\eta_{\theta}\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t})$, where $\mathcal{L}_{\mathcal{S}}$ and $\mathcal{L}_{\mathcal{T}}$ denote the task losses on the synthetic and original graphs, respectively, and $D(\cdot,\cdot)$ is a distance metric between gradient vectors.
Although traditional gradient matching methods succeed in unimodal settings, they struggle on multimodal graphs due to gradient conflicts and structural noise amplification. Our analysis Fig. 1 shows that standard gradient matching compresses inter-modal differences during training, leading to modality collapse—where multimodal information degenerates into a unimodal representation. Moreover, GNN message passing amplifies gradient noise, correlating update directions across nodes and accelerating feature concentration (higher CR). In the following sections, we will analyze these two issues in detail.
@@ 2.3. Inter-Modality Gradient Conflict
In the gradient matching process for multimodal graph condensation, directional conflicts occur between the feature gradients of different modalities within a node, the root of which can be traced back to the inherent opposition of modal gradients in the parameter space. The parameter gradient of the original graph, $\nabla_{\theta}\mathcal{L}_{T}$, can be decomposed into independent components for the text and image modalities, $v_{t}$ and $v_{i}$. When the optimal update directions of the two modalities are inconsistent, the angle between them satisfies $\psi>\pi/2$, resulting in an inner product $\langle v_{t},v_{i}\rangle<0$, thereby forming a modal conflict at the parameter level.
Gradient matching aims to make the synthetic gradient $\nabla_{\theta}\mathcal{L}_{S}$ approximate this target, i.e., $\nabla_{\theta}\mathcal{L}_{S}\approx\alpha\nabla_{\theta}\mathcal{L}_{T}$. The synthetic gradient is connected to the node feature gradients via the Jacobian matrices $A_{i}$ and $B_{i}$.
Specifically, let the text feature of node $i$ be $x_{i,\text{text}}$ and its image feature be $x_{i,\text{image}}$. Their corresponding feature gradients are $g_{i,\text{text}}=\frac{\partial\mathcal{L}_{S}}{\partial x_{i,\text{text}}}$ and $g_{i,\text{image}}=\frac{\partial\mathcal{L}_{S}}{\partial x_{i,\text{image}}}$, respectively. The Jacobian matrices $A_{i}=\frac{\partial}{\partial x_{i,\text{text}}}\nabla_{\theta}\mathcal{L}_{S}\in\mathbb{R}^{|\theta|\times d_{t}}$ and $B_{i}=\frac{\partial}{\partial x_{i,\text{image}}}\nabla_{\theta}\mathcal{L}_{S}\in\mathbb{R}^{|\theta|\times d_{v}}$ represent the partial derivatives of the parameter gradient with respect to the node’s text and image features, respectively. They linearly map the node feature gradients into the parameter gradient space. Therefore, the synthetic gradient can be expressed as:
(3) $\nabla_{\theta}\mathcal{L}_{S}=\sum_{i=1}^{N^{\prime}}\left(A_{i}\,g_{i,\text{text}}+B_{i}\,g_{i,\text{image}}\right)$
Due to parameter sharing and feature fusion, the column spaces of $A_{i}$ and $B_{i}$ often overlap (modal mixing). To satisfy the matching condition $A_{i}\,g_{i,\text{text}}+B_{i}\,g_{i,\text{image}}\approx\alpha(v_{t}+v_{i})$, the optimization process must reconcile the opposing directions of $v_{t}$ and $v_{i}$ within the overlapping subspace. This forces the projections of the feature gradients $g_{i,\text{text}}$ and $g_{i,\text{image}}$ onto this subspace to become opposite in direction, thereby triggering intra-node feature gradient conflict:
(4) $\langle g_{i,\text{text}},g_{i,\text{image}}\rangle<0$
Therefore, intra-node modal feature gradient conflict is an inevitable outcome when gradient matching seeks to fit a multimodal target that is inherently conflicted, under a model structure with modal mixing. The key to addressing this issue lies in gradient decoupling. As shown in Fig. 1(b), decoupling such conflict effectively preserves inter-modal feature diversity and prevents collapse.
@@ 2.4. Structural Amplification of Gradient Noise
The meta-gradient calculation in gradient matching involves backpropagation through the GNN architecture, which amplifies the noise in modal gradients. In the outer loop, the update direction of the synthetic graph is given by:
(5) $\Delta\mathbf{X}_{\mathcal{S}}=-\eta\nabla_{\mathbf{X}_{S}}D(\nabla_{\theta}\mathcal{L}_{\mathcal{S}},\nabla_{\theta}\mathcal{L}_{T})$
@@
This meta-gradient is computed via the chain rule, which requires backpropagation through the message-passing layers of the GNN. Specifically, during meta-gradient computation, the gradient of a node’s features is influenced by the gradients of its neighbors, causing gradients to propagate backward along the graph edges:
(6) $\nabla_{h_{i}^{(k-1)}}\mathcal{L}=\sum_{j\mid i\in\mathcal{N}(j)}\frac{\partial\mathcal{L}}{\partial h_{j}^{(k)}}\frac{\partial h_{j}^{(k)}}{\partial h_{i}^{(k-1)}}.$
@@
This dynamic transforms the computational graph of the GNN into a medium for noise propagation. The gradient conflicts at local nodes are not contained; instead, they are amplified and propagated throughout the graph structure. Let the gradient noise of node representations at the $l$-th layer be denoted by $\mathbf{R}^{(l)}$. Then the noise propagation in backpropagation can be formalized as:
(7) $\mathbf{R}^{(l-1)}=\mathbf{A}_{S}^{\top}\mathbf{R}^{(l)}\mathbf{W}^{(l)\top}+\mathbf{E}^{(l)},$
where $\mathbf{A}_{S}$ is the adjacency matrix of the synthetic graph, $\mathbf{W}^{(l)}$ is the weight matrix of the $l$-th layer, and $\mathbf{E}^{(l)}$ is the inherent modal mixing noise at that layer. This results in a highly non-smooth gradient field over the synthetic graph, where the optimization directions of adjacent nodes can differ significantly.
To quantify this effect, we define a noise amplification operator on the graph, $\mathcal{P}:\mathbb{R}^{N^{\prime}\times d}\to\mathbb{R}^{N^{\prime}\times d}$, as $\mathcal{P}(\mathbf{R})=\mathbf{L}_{S}\mathbf{R}$, where $\mathbf{L}_{S}$ is the Laplacian matrix of the synthetic graph and $\mathbf{R}$ is the node gradient noise matrix. This operator captures the amplification effect of the graph structure on noise. Based on this, we obtain the following theorem:
Theorem 2.1.
Under gradient matching optimization, the amplification factor of modal mixing noise propagated through the graph structure is bounded by the Dirichlet energy of the gradient field:
(8) $\|\mathbf{L}_{S}\mathbf{R}\|_{F}^{2}\leq\lambda_{\max}(\mathbf{L}_{S})\cdot E(\mathbf{R}),$
where $E(\mathbf{R})=\operatorname{tr}(\mathbf{R}^{\top}\mathbf{L}_{S}\mathbf{R})$ is the Dirichlet energy of the gradient field, and $\lambda_{\max}(\mathbf{L}_{S})$ is the maximum eigenvalue of the Laplacian matrix.
The proof of Theorem 2.1 is provided in Appendix A.5. Theorem 2.1 shows that the graph structure systematically amplifies gradient noise through its spectral properties, thereby exacerbating optimization instability. In particular, a high Dirichlet energy corresponds to a chaotic gradient field, which the structural amplification mechanism sustains, resulting in a rugged optimization landscape and destabilized meta-gradient. This persistent instability directly hinders the condensation objective.
We further analyze the impact of Dirichlet energy $E(\mathbf{R})$ on gradient field stability and condensation performance through experiments. The results confirm a clear negative correlation Fig. 2(a): lower $E(\mathbf{R})$ consistently leads to better condensed graph performance. Thus, suppressing this energy field stabilizes the gradient field and enhances condensation performance, as shown in Fig. 2(b).
Figure 3. Performance and Dirichlet energy [^2] Dirichlet energy (Zhou et al., 2021) is a key metric that measures the smoothness of graph signals or embeddings over the graph topology.analysis. (a): Relationship between node classification accuracy and $E(\mathbf{R})$ during a condensation process; (b): Evolutionary trends of $E(\mathbf{R})$ for GCond and SR-GM. For clarity, $E(\mathbf{R})$ is normalized and log-transformed.